OpenStudy (anonymous):

ok so im working with imaginary numbers and in this particular equation i have to solve for b. (3+2i)+(2+bi)=5-4i

jhonyy9 (jhonyy9):

how you make calcules with imaginary numbers ?

jhonyy9 (jhonyy9):

3+2i +2+bi =

jhonyy9 (jhonyy9):

just assum the same terms

OpenStudy (anonymous):

im sorry i dont understand what you mean

jhonyy9 (jhonyy9):

3+2 = ? and from 2i and bi what can you factorizing out ?

OpenStudy (anonymous):

ok im failing this class so your going to have to break it down further than that luis

jhonyy9 (jhonyy9):

yes i like to learn you calcule it step by step if you like it the same too ?

OpenStudy (shubhamsrg):

just treat i as any other constant.. 3+ 2i + 2 + bi = (3+2) + i(2+b) = 5 -4i for a+ ib = x+ iy ..we always have a= x and b=y that is 2 complex no.s are equal if and only if both the real and imaginary parts are equal.. so equate the imaginary parts on both sides and solve for b..

jhonyy9 (jhonyy9):

so now you have understood it ?

OpenStudy (anonymous):

yes i get it now thank you everyone

jhonyy9 (jhonyy9):

ok good luck bye